OpenStudy (anonymous):

Will someone help me with circles. I'm just not catching on with it and I have 2 assignments to do with them..

OpenStudy (phi):

what is the question?

OpenStudy (anonymous):

number 1 right now but I need help with it all I just don't understand, im sure if you help me with 1 ill get a few others

OpenStudy (anonymous):

ignore the answer to #1 im just trying to try stuff to help but its not helping that I don't have multiple choice

OpenStudy (phi):

I think they want you to know (remember) that the measure of an arc is equal to the measure of the central angle. (the central angle is the same as the measure of the arc and vice versa)

OpenStudy (anonymous):

so that angle will alao be 102 degrees?

OpenStudy (phi):

yes

OpenStudy (anonymous):

also*

OpenStudy (anonymous):

thank you

OpenStudy (phi):

for Q2, they want the smaller central angle, but give you the "bigger angle" remember all the way around is 360º

OpenStudy (anonymous):

okay, so I would take 360-302=58. and then 58 degrees is what that angle would be?

OpenStudy (phi):

yes

OpenStudy (phi):

For Q3, you should know (memorize) an inscribed angle is ½ of the central angle (or, the same thing, ½ of its arc)

OpenStudy (anonymous):

im working that one out and ill tell you what I got and see if im doing it right

OpenStudy (anonymous):

okay, I did 360-(x+45) then got it down to 315-x then I set the angle to it x=315-x and got 157.5 and then took half of that and got 78.75

OpenStudy (phi):

pretty close. arc AB= 315 - x the inscribed angle ACB is ½ of arc AB in other words <ACB = ½ (315-x) we know <ACB is x, so this is the same as x = ½ (315-x) notice we have to use the ½ idea first (not at the end)

OpenStudy (anonymous):

ohh. im still stuck. Would I still get 78.75 putting 1/2 first

OpenStudy (phi):

did you start with \[ x = \frac{1}{2}(315-x) \]?

OpenStudy (phi):

the first step could be, multiply both sides by 2 what do you get ?

OpenStudy (anonymous):

2x=630-2x???

OpenStudy (phi):

like this \[ 2 \cdot x =2 \cdot \frac{1}{2}(315-x)\] the 2 * ½ is 1 so this becomes \[ 2x = 315 -x \] now add +x to both sides.

OpenStudy (anonymous):

3x=315 105

OpenStudy (phi):

yes. See if you can make sense of what we did to solve this one (which was a bit tricky)

OpenStudy (anonymous):

I understand but then I don't. But I had an example in the lesson and wrote down step by step but that last part got me

OpenStudy (phi):

The first trick part is finding arc AB is 315-x you did that. then you need to know that angle ACB is ½ of arc AB angle ACB = ½ (315 -x) they tell you angle ACB is equal to x, so you get the equation \[ x = \frac{1}{2} (315-x) \] to solve you need to know algebra.

OpenStudy (anonymous):

okay thanks I understand that a lot better. The only question I have about the next one is how would I solve it without the major arc. I thought I had to take the major arc minus 360 then divide that by 2 and set that answer equal to each equation

OpenStudy (phi):

for Q4, angle x is how much of arc WZ ?

OpenStudy (anonymous):

half?

OpenStudy (phi):

yes. now what about angle y? what arc is it ½ of ?

OpenStudy (anonymous):

WZ

OpenStudy (phi):

in other words angle x = WZ/2 and angle y = WZ/2 that means angle x equals angle y

OpenStudy (anonymous):

so x+2=3x-40

OpenStudy (phi):

yes. now it's algebra ....

OpenStudy (anonymous):

thank you im going to work it out

OpenStudy (anonymous):

I got 42?

OpenStudy (phi):

you need to practice your algebra. x+2=3x-40 add -x to both sides: -x+x+2 = 3x - x - 40 simplify: 2 = 2x -40 can you finish?

OpenStudy (anonymous):

yes, equations are usually my easiest thing. I think I just need to focus more. I think i got the rest of the assignment. If I have any more questions i'll come to you. Thank you! you helped a lot:)

OpenStudy (anonymous):

@phi I ended up getting a 100 on this assignment, thank you